# Testing Dark Matter with Generative Models for Extragalactic Stellar Streams

### v1

@@ 3.1 Stream Generation and Model Parameters
Table 2: Priors on model parameters
@@
3.2 Mock Observations
[…]
@@ 3.3 Cost function
Our cost function is the Kullback–Leibler (KL) divergence (Kullback & Leibler, 1951) between the distribution of control points ($p_{\rm data}$) and the model distribution ($p_{\rm model}$). Statistically, one can interpret the KL divergence as an average log-likelihood ratio:
@@
[…]
3.4 Sampling
From Eq. 5, the optimal $\theta$ will minimize the KL divergence. We therefore seek to maximize the function $-\rm{KL}(\theta)$. We accomplish this using nested sampling, implemented in the nautilus python package (Lange, 2023). Unlike standard nested sampling techniques which involve a rejection step, nautilus employs importance sampling, allowing every proposal sample to contribute to evaluating the posterior distribution. Additionally, the proposal distribution is adaptively learned using multilayer perceptron neural networks, which significantly accelerates the inference process.
While these features of the nautilus code enhance the efficiency of the sampling procedure, the most substantial speedup comes from the vectorized nature of our simulator. Specifically, we use the vmap (vectorized map) functionality in the JAX framework to simultaneously simulate 500 different parameter sets, and thus stream realizations. This parallelization of likelihood evaluations enables much faster exploration of the parameter space, reducing the runtime of our method from over 10 hours to, at most, a few hours. Further speedups may be achieved in future iterations, as discussed in §5.2.
3.4.1 Drawing Contours
[…]
@@ 3.4.1 Drawing Contours
In Appendix B, we validate that Eq. 9 predicts measured values of $\Lambda$ (Eq. 6) under the conditions of normality. In §5, we will illustrate that even for more complicated distributions, Eq. 9 provides a means to enclose regions of parameter space containing samples that match the input data.
[…]
4 Results
4.1 Parameter constraints
We apply our method to the two streams introduced in §2. Our sampler has 10 free parameters and 5 fixed parameters. See Table 2 for the prior ranges for both the inner and outer stream.
[…]
We also find that the method naturally produces limits on the halo mass ($M_{\rm halo}$). These constraints arise because the mock streams have specific widths, which enters into the stream model through the tidal radius: $r_{t}\propto\left(\frac{M_{\rm prog}}{M_{\rm enc}}\right)^{1/3}$. Thus selecting a range for the progenitor mass while keeping the stream width fixed will in turn limit the halo mass. For an otherwise fixed progenitor, increasing the halo mass will typically decrease the width of streams. This sets an upper bound on $M_{\rm halo}$ (Erkal et al., 2016), since too massive of a halo will produce a stream narrower than the observations. We also reject unbound progenitors, which further limits the maximum mass of our halo model. Additional constraints come from the observed length of the stream. Streams shorter than the observed length are disfavored. This places a lower limit on the halo mass: if the mass is too small, the resulting stream will be typically shorter at a fixed distance compared to the data. Note, however, that there is also a degeneracy with the dynamical age, $t_{age}$ of the stream (see §4.2) and $M_{\rm prog}$. The stream morphology is not highly sensitive to the scale radius, $r_{s}$, or to the stream age as indicated by the flat posterior distributions. Note that while mass bound can be obtained for the simple particle spray method we have used here, real tidal stripping is more complicated and mass constraints should be viewed as potentially highly model dependent. We defer an exploration of the robustness of mass constraints to future work.
To demonstrate that the sampler successfully identifies orbital solutions and halo profiles consistent with the mock observations, we present in Fig. 6 the maximum likelihood forward models for the inner stream (top left) and the outer stream (bottom left). Model streams are compared to the input data (gray histograms), and show excellent agreement. The middle panels of Fig. 6 display stream models corresponding to the blue and red points (consistent within 68%) and green points (inconsistent) from Figures 5 and 12. The blue model corresponds to a scenario in which the progenitor is located behind the host galaxy at the present day (i.e., negative $y_{prog}$), while the red model has the progenitor in front of the host galaxy (i.e., positive $y_{prog}$). Both of these models closely match the input data. In contrast, the green models, shown in the rightmost column, provide a poor fit to the input data as expected. Note we have illustrated one degeneracy here (line-of-sight distance), additional degeneracies will be discussed in §4.2.
@@ 4.1 Parameter constraints
In the lower panel of Fig. 7 we plot the fractional uncertainty in the inferred density relative to the true value. The inner stream provides a constraint that is approximately five times tighter in the inner regions of the halo compared to the outer stream. Both streams yield similar constraints on the outer slope, though the outer stream performs slightly better. The minimum uncertainty for the inner stream occurs around $15~\rm{kpc}$, while for the outer stream, it occurs around $26~\rm{kpc}$. The shaded gray region in the bottom panel denotes the prior range for the scale radius, $r_{s}$. We highlight this region because the minimum uncertainty on the density profile is expected to fall in this range, because the scale-radius marks the transition region between the inner density power-law and the outer density power-law. Still, even within the prior range, the radial location of the outer stream’s best constraint is at a distance approximately twice of the inner stream’s best constraint. This supports the intuition that each stream constrains the region of the potential local to its orbit (e.g., Bonaca & Hogg 2018; Nibauer et al. 2022), with closer-in streams providing more information on the inner regions of the halo.
Figure 7: Top: Constraint on the radial density profile from the inner (blue), outer (red), and combined (black dashed curves) stream samples. The true profile is NFW (black line), and the gray region shows the allowed density profiles from our priors. Both streams show comparable constraints on the outer slope, as their orbits are typically a factor of 2×2$\times$2 × the halo scale radius. However the inner stream places tighter constraints on the inner slope, since its orbit reaches the scale radius of the halo. Bottom: fractional uncertainty on the radial profile of the halo as a function of distance. The inner stream’s errorbar is a factor of 5$\approx$ 5$\times$ 5 × tighter at small radii, while the outer stream offers marginally tighter constraints on the outer density profile compared to the inner stream. The gray shaded region shows the prior range on the scale radius, $r_{s}$.
@@
4.2 Degeneracies
[…]
In the second panel of Fig. 8 we highlight the degeneracy between the progenitor mass, $M_{\rm prog}$ and the dynamical age, $t_{age}$. For a more massive progenitor, the particle-spray method produces a larger scatter in velocity as stars escape the Lagrange points (see §3), which leads to more rapid phase mixing thereby producing longer streams. This means that the same length stream can be achieved by increasing the progenitor mass, while decreasing the dynamical age. The two samples illustrated from the high probability regions are extremely similar in radial velocity and morphology.
In the third panel of Fig. 8, we show the inner slope, $\gamma$, halo mass degeneracy, and in the fourth panel we show the outer slope, $\beta$, halo mass degeneracy. For higher $\gamma$, we see a degeneracy that correlates with lower halo mass. This confirms our intuition, as larger $\gamma$, i.e. a steeper inner density slope, leads to more rapid phase mixing and thus longer streams (see §2). To produce a stream with the same length, we in turn need to lower the halo mass. In the fourth panel, we see that a steeper outer slope (higher $\beta$) leads to a degeneracy correlating with higher halo mass. This also confirms our intuition that higher $\beta$ truncates the intermediate-outer regions of the density, so a higher overall mass amplitude is needed to match the input data (see Fig. 1). The two streams randomly sampled within the 68% region of the likelihood surface do not show differences in morphology or radial velocity (see top panels).
@@ 4.2 Degeneracies
4.3 Cored density profiles
We have showed that our method can recover the inner and outer slope of a NFW profile. Here we test whether our approach can also recover the true potential parameters if the stream is evolving in a dark matter halo with a cored density profile. Cored profiles are particularly interesting for dark matter physics, since they are expected outcomes in self-interacting dark matter (Spergel & Steinhardt, 2000; Bullock & Boylan-Kolchin, 2017). As the outer stream was not as sensitive to the inner slope ($\gamma$) of the potential (see Fig. 2), we attempt to recover $\gamma$ with the inner stream. To run our test, we simulate the inner stream in a cored potential with $\gamma=0.2$ and $\beta=3$.
[…]
@@ 4.3 Cored density profiles
5 Discussion
In this section, we discuss how our method compares to other approaches for modeling extragalactic tidal streams (§5.1), and discuss the limitations of our method (§5.2).
5.1 Comparison to Prior Work
@@
Walder et al. (2024) developed a method to fit orbits to extragalactic stellar streams in a power-law potential of the form $\Phi\propto r^{-\gamma}$. They fit orbits to mock-streams launched from apocenter as a function of different inclination angles. We have reproduced their findings that it is possible to infer the inner slope of the potential (or density) from the turning points of a stream’s track. We have improved upon their method in two main regards. First, we do not fit orbits to the stream’s track, but use a full generative stream model to compare to the observations. Because streams can be significantly misaligned with their progenitor’s orbit, orbit fitting can lead to bias (Sanders & Binney, 2013). Additionally, our approach includes stream width and length information as constraints on the potential. We find that in combination, these can be used to set mass bounds on the halo. Second, our model has more flexibility to model the underlying stream and dark matter halo. In particular, we model the progenitor’s mass and age, while also allowing for variations in an inner and outer halo density slope. We find that multiple streams can be used to stitch together the full radial profile. The disadvantage of our approach is that it is likely slower than the method from Walder et al. (2024), because we compute 100010001000 orbits per stream realization. It is worth considering whether a hybrid technique that uses less orbit integrations while also retaining the flexibility we highlight above could lead to further improvements in speed (e.g., the streakline approach Küpper et al. 2012).
5.2 Limitations
@@
For dwarf galaxies, however, much of the stripping occurs near pericenter (see e.g. Bonaca & Price-Whelan, 2025). In Pearson et al. (2022b), they included a Jacobi radius stripping criterion for their dwarf progenitor, which ensured that there was no tidal stripping if the Jacobi radius was much larger than the extent of the progenitor. This in turn led to preferred stripping near pericenter, where the Jacobi radius is smaller. Adding a Jacobi radius criterion to our method would only make it more robust and able to reject more orbits. We can relax the assumption of uniform stripping similar to that of Pearson et al. (2022b) in future work.
Progenitor location: Throughout this paper we fix the present day on-sky position of the progenitor ($x,z$). In observations we will often not have access to the progenitor’s sky position, which introduces some uncertainty to our method (e.g., Mart´ınez-Delgado et al., 2023; Miró-Carretero et al., 2024a). In future work, we can overcome this by marginalizing the progenitor’s location over the stream’s track. This would add a parameter (the angular location of the progenitor along the track), and likely introduce a new set of degeneracies. The method will therefore be more computationally expensive, as we would have to explore many more orbits. However, this additional complexity will most clearly impact our inference of the progenitor properties, but to a lesser extent the radial profile of the host.
@@
Mock observing the streams: In this paper, we applied an “observational” threshold criteria (see §3), but otherwise assumed perfect data. In observations, we likely only observe the densest parts of streams, which will play into the estimated width of each stream from the control points. This could be addressed by adding uncertainties or scatter to our control points representing each stream (see Fig. 4). We do, however, ensure that streams longer than the input data are not disfavored in our sampling.
@@
In halos where two streams are observed, it can be difficult to disentangle whether the two streams are indeed distinct or from the same progenitor (see e.g., Fielder et al., 2025, E. Bell et al., in prep.). Miro-Carretero et al. (2024) found that if we reach a surface brightness limit of 31 mag arcsec-2 stellar tidal streams should be detected around 50% of surveyed hosts, however the occurrence of two streams around one host galaxy might be low. The Roman wide field of view combined with its depth will be ideal for such studies. We also note that streams at smaller galactocentric radii, such as the inner stream, could be more difficult to detect due to a lower surface brightness contrast to the stellar halo. While the detection of multiple dwarf galaxy streams around one host galaxy might be rare and difficult to disentangle, our method can also be used on extragalactic globular cluster streams, predicted to be discovered by Roman (Pearson et al., 2019, 2022a), or on dwarf streams around dwarf galaxies (e.g. Mart´ınez-Delgado et al., 2012; Kado-Fong et al., 2018; Carlin et al., 2019; Fielder et al., 2025).
@@
Time-dependence: Throughout this paper, we have assumed that the host potential and progenitor mass are static. From the Milky Way, we know that time dependence induced from accreted satellites (e.g., Erkal et al., 2019; Shipp et al., 2021; Vasiliev et al., 2021; Arora et al., 2022; Dillamore et al., 2022; Lilleengen et al., 2023; Nibauer et al., 2024; Brooks et al., 2025) and from the overall mass growth of host halos (Buist & Helmi, 2015) can affect the orbits and morphologies of stellar streams. However, depending on the location of the stream in its host halo, the recent merger history, and the dynamical age of the streams, some streams may be less sensitive to time dependence (Nibauer et al., 2022; Brooks et al., 2025). Buist & Helmi (2015) showed that for a smooth halo growth that well reproduced the average mass accretion history of the Milky-Way-sized halos, time evolution led to an angular difference in apocenters at the level of $\approx 10~\deg$, but only for sufficiently long streams which wrapped around their host more than once. In an upcoming work, we test our method on cosmologically evolved halos, where we can determine the bias induced by realistic time-dependence.
Spherical halo assumption: We have assumed a spherical potential to represent our host halos. $\Lambda$CDM predicts that dark matter halos are often triaxial (Frenk et al., 1988; Dubinski & Carlberg, 1991; Warren et al., 1992; Jing & Suto, 2002; Vera-Ciro et al., 2011; Arora et al., 2025), though more spherical profiles in the inner regions are possible depending on the mass of the baryons (Kazantzidis et al., 2010). Non-spherical halos will allow for orbital plane precession (Erkal et al., 2016; Nibauer et al., 2024), which is in turn degenerate with the precession of apocenters in halos with different density slopes. Nibauer et al. (2023) showed that the curvature of stellar streams can be used to infer the flattening of halos, and their (mis)alignment with a disk component. It is possible to use the method of Nibauer et al. (2023) to rapidly determine which streams prefer spherical geometries, and target those candidates using the method presented here. Alternatively, it is possible to include flattening in our sampling procedure, though the runtime of our generative method is on the order of hours, while the method of Nibauer et al. (2023) works in seconds. In future work, we intend to use the curvature method to inform prior distributions for X-Stream.

### v2

@@ III.1 Stream Generation and Model Parameters
Table 2: Priors on model parametersfootnotetext:
@@
III.2 Mock Observations
We create mock observations of the inner and outer streams which we use as the input for our method. In Fig. 4 (left) we show the particle positions of the inner and outer streams (same the left panels of Fig. 2). For current and future stream imaging data (e.g. from Euclid, Roman, or LSST), we have access to the on-sky morphology of stellar streams. However, the faintest parts of the streams will be difficult to detect (see e.g., Bullock and Johnston, 2005; Miró-Carretero et al., 2025). Since we are working with particle data from simulations, we construct histogram representations of the two streams to emulate mock observations (Fig. 4, middle panel). We then place representative control points in each bin that contains a few stars. For the inner stream, we use 30 bins in $x\in[-50,50]~\rm{kpc}$ and $z\in[-60,50]~\rm{kpc}$. The outer stream spans roughly twice the area, so we double the bin size, using $x\in[100,80]$ and $z\in[20,180]$. We apply a threshold mask by setting the count of bins to zero if they contain fewer than $0.2\%$ of the total number particles in each stream’s tidal tail. All model streams consist of 1000 particles. However, for the mock data, we increase the resolution: the inner stream contains 2000 particles, and the outer stream contains 8000 particles. The higher particle count is used solely to automate the selection of control points. With too few particles, the control points become sparse and unrepresentative of the actual stream structure, and their distribution becomes highly sensitive to bin size. This limitation does not affect our sampling procedure, as binning is used only to generate mock observations–not for comparing models to data.
@@ III.3 Cost function
Our cost function is the Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951) between the distribution of control points ($p_{\rm data}$) and the model distribution ($p_{\rm model}$). Statistically, one can interpret the KL divergence as an average log-likelihood ratio:
@@
where $\theta$ are model parameters, $p_{\rm data}$ is the prior probability of control points ($1/N$), and the average is over the control points ($x_{i},z_{i}$). We use nested sampling to generate samples from Eq. 5. Details of the nested sampling procedure will be discussed in §III.4. Once we generate a distribution of parameters, $\theta$, we must then draw $68\%$ and $95\%$ high confidence regions using the procedure defined in §III.4.1.
III.4 Sampling
From Eq. 5, the optimal $\theta$ will minimize the KL divergence. We therefore seek to maximize the function $\rm{KL}(\theta)$. We accomplish this using nested sampling, implemented in the nautilus python package (Lange, 2023). Unlike standard nested sampling techniques which involve a rejection step, nautilus employs importance sampling, allowing every proposal sample to contribute to evaluating the posterior distribution. Additionally, the proposal distribution is adaptively learned using multilayer perceptron neural networks, which significantly accelerates the inference process.
While these features of the nautilus code enhance the efficiency of the sampling procedure, the most substantial speedup comes from the vectorized nature of our simulator. Specifically, we use the vmap (vectorized map) functionality in the JAX framework to simultaneously simulate 500 different parameter sets, and thus stream realizations. This parallelization of likelihood evaluations enables much faster exploration of the parameter space, reducing the runtime of our method from over 10 hours to, at most, a few hours. Further speedups may be achieved in future iterations, as discussed in §V.2.
III.4.1 Drawing Contours
Having generated samples from $\rm{KL}(\theta)$, we must now decide which samples belong to regions of $68\%$ and $95\%$ confidence. We will make a number of simplifying assumptions in converting our samples of Eq. 5 to posterior estimates, and validate the quality of the posteriors in §IV. Note that we take a similar statistical approach to Sanderson et al. (2015), who used the KL divergence as a test-statistic and then determined regions of 68 and 95% confidence as a post-processing step.
@@ III.4.1 Drawing Contours
From Eq. 9 we can find the value of $\Delta\sigma$ corresponding to the $\Lambda$ value calculated for each model (Eq. 6). That is, we solve for the inverse $\Delta\sigma=F^{-1}\left({\Lambda}\right)$, where $F^{-1}$ is the inverse of Eq. 9. We fit for the inverse using spline interpolation. Because $\Delta\sigma$ represents a deviation from the best fit, it is effectively a $z$-score. We therefore can compute the probability of a $\Delta\sigma$ value using the standard normal, $p(\Delta\sigma)\propto\exp\left[-\frac{1}{2}(\Delta\sigma)^{2}\right]$. This allows us to determine a weight for each sample, which is then used to enclose regions of $68$ and $95\%$ confidence.
In Appendix B, we validate that Eq. 9 predicts measured values of $\Lambda$ (Eq. 6) under the conditions of normality. In §V, we will illustrate that even for more complicated distributions, Eq. 9 provides a means to enclose regions of parameter space containing samples that match the input data.
We note that because we sample the negative KL-divergence, the density of samples can be different from the true posterior (i.e., the average in Eq. 5 can suppress sharp peaks in $p_{\rm model}$). To correct for this, posterior constraints are derived by constructing bins in parameter space and taking the average weight of the samples in each bin. Contours are then drawn using the average weight in each bin. This approach ensures that contours are drawn according to the weights rather than the sampling density of the tempered likelihood.
Figure 5: Constraints on the 10 free parameters for the inner stream. The black lines show the true parameters. The red and blue points represent two randomly selected good fits within the 68% credible region, while the green point corresponds to a poor fit. We visualize the streams corresponding to these points in Fig. 6. The sampler recovers all true parameters within the $68\%$ region, and places strong limits on the inner slope of the density profile, $\gamma$. There are minimal 1D marginal constraints on $v$, $r_{s}$, $t_{age}$, though these quantities are degenerate with $M_{\rm halo}$, and $t_{\rm age}$ is degenerate with $M_{\rm prog}$. There is information on the line-of-sight distance, $y_{prog}$, and velocity directions, $\theta,\phi$.
IV Results
IV.1 Parameter constraints
We apply our method to the two streams introduced in §II. Our sampler has 10 free parameters and 5 fixed parameters. See Table 2 for the prior ranges for both the inner and outer stream.
@@ IV.1 Parameter constraints
We also find that the method naturally produces limits on the halo mass ($M_{\rm halo}$). These constraints arise because the mock streams have specific widths, which enters into the stream model through the tidal radius: $r_{t}\propto\left(\frac{M_{\rm prog}}{M_{\rm enc}}\right)^{1/3}$. Thus selecting a range for the progenitor mass while keeping the stream width fixed will in turn limit the halo mass. For an otherwise fixed progenitor, increasing the halo mass will typically decrease the width of streams. This sets an upper bound on $M_{\rm halo}$ (Erkal et al., 2016), since too massive of a halo will produce a stream narrower than the observations. We also reject unbound progenitors, which further limits the maximum mass of our halo model. Additional constraints come from the observed length of the stream. Streams shorter than the observed length are disfavored. This places a lower limit on the halo mass: if the mass is too small, the resulting stream will be typically shorter at a fixed distance compared to the data. Note, however, that there is also a degeneracy with the dynamical age, $t_{age}$ of the stream (see §IV.2) and $M_{\rm prog}$. The stream morphology is not highly sensitive to the scale radius, $r_{s}$, or to the stream age as indicated by the flat posterior distributions. Note that while mass bound can be obtained for the simple particle spray method we have used here, real tidal stripping is more complicated and mass constraints should be viewed as potentially highly model dependent. We defer an exploration of the robustness of mass constraints to future work.
To demonstrate that the sampler successfully identifies orbital solutions and halo profiles consistent with the mock observations, we present in Fig. 6 the maximum likelihood forward models for the inner stream (top left) and the outer stream (bottom left). Model streams are compared to the input data (gray histograms), and show excellent agreement. The middle panels of Fig. 6 display stream models corresponding to the blue and red points (typical points from the reweighted samples) and green points (inconsistent) from Figures 5 and 12. We also plot other examples of inconsistent streams that are not supported by our sampling in Fig. 6 (orange and brown models). The blue model corresponds to a scenario in which the progenitor is located behind the host galaxy at the present day (i.e., negative $y_{prog}$), while the red model has the progenitor in front of the host galaxy (i.e., positive $y_{prog}$). Both of these models closely match the input data. In contrast, the green, brown, and orange models shown in the rightmost column provide a poor fit to the input data as expected. Note we have illustrated one degeneracy here (line-of-sight distance), additional degeneracies will be discussed in §IV.2.
@@
In the lower panel of Fig. 7 we plot the fractional uncertainty in the inferred density relative to the true value. The inner stream provides a constraint that is approximately five times tighter in the inner regions of the halo compared to the outer stream. Both streams yield similar constraints on the outer slope, though the outer stream performs slightly better. The minimum uncertainty for the inner stream occurs around $15~\rm{kpc}$, while for the outer stream, it occurs around $26~\rm{kpc}$. The shaded gray region in the bottom panel denotes the prior range for the scale radius, $r_{s}$. We highlight this region because the minimum uncertainty on the density profile is expected to fall in this range, because the scale-radius marks the transition region between the inner density power-law and the outer density power-law. Still, even within the prior range, the radial location of the outer stream’s best constraint is at a distance approximately twice of the inner stream’s best constraint. This supports the intuition that each stream constrains the region of the potential local to its orbit (e.g., Bonaca and Hogg 2018; Nibauer et al. 2022), with closer-in streams providing more information on the inner regions of the halo.
Figure 7: Top: Constraint on the radial density profile from the inner (blue), outer (red), and combined (black dashed curves) stream samples. The true profile is NFW (black line), and the gray region shows the allowed density profiles from our priors. Both streams show comparable constraints on the outer slope, as their orbits are typically a factor of $2\times$ the halo scale radius. However the inner stream places tighter constraints on the inner slope, since its orbit reaches the scale radius of the halo. Bottom: fractional uncertainty on the radial profile of the halo as a function of distance. The inner stream’s errorbar is a factor of $\approx 5\times$ tighter at small radii, while the outer stream offers marginally tighter constraints on the outer density profile compared to the inner stream. The gray shaded region shows the prior range on the scale radius, $r_{s}$.
@@
IV.2 Degeneracies
In Fig. 8, we illustrate some of the degeneracies for the inner stream. The panels show examples of two streams sampled within the $68\%$ region, colored by line-of-sight velocity. In the left panel we show the angle $\phi$, which controls how much of the progenitor’s velocity is along the line-of-sight, versus $y_{\rm{prog}}$, the progenitor’s line-of-sight location. The bimodality in $\phi$ shows the leading or trailing tail can be swapped without impacting the fit, as expected, since there is not a clear “S”-shape for either stream in our mock-observations. The modes are separated by roughly $\pi$ radians, which is also the case for the outer stream (see Fig. 12). We plot two streams from each of the peaks in the likelihood surface (see Fig. 8 top left panels), and recover the Pearson et al. (2022b) result (see their Fig. 10), which shows a flipped radial velocity sign for a stream which has its progenitor in front of or behind the host in line-of-sight distance. Red indicates positive radial velocity (moving away from the observer), and blue indicates negative radial velocity (moving towards the observer).
In the second panel of Fig. 8 we highlight the degeneracy between the progenitor mass, $M_{\rm prog}$ and the dynamical age, $t_{age}$. For a more massive progenitor, the particle-spray method produces a larger scatter in velocity as stars escape the Lagrange points (see §III), which leads to more rapid phase mixing thereby producing longer streams. This means that the same length stream can be achieved by increasing the progenitor mass, while decreasing the dynamical age. The two samples illustrated from the high probability regions are extremely similar in radial velocity and morphology.
In the third panel of Fig. 8, we show the inner slope, $\gamma$, halo mass degeneracy, and in the fourth panel we show the outer slope, $\beta$, halo mass degeneracy. For higher $\gamma$, we see a degeneracy that correlates with lower halo mass. This confirms our intuition, as larger $\gamma$, i.e. a steeper inner density slope, leads to more rapid phase mixing and thus longer streams (see §II). To produce a stream with the same length, we in turn need to lower the halo mass. In the fourth panel, we see that a steeper outer slope (higher $\beta$) leads to a degeneracy correlating with higher halo mass. This also confirms our intuition that higher $\beta$ truncates the intermediate-outer regions of the density, so a higher overall mass amplitude is needed to match the input data (see Fig. 1). The two streams randomly sampled within the 68% region of the likelihood surface do not show differences in morphology or radial velocity (see top panels).
@@ IV.2 Degeneracies
IV.3 Cored density profiles
We have showed that our method can recover the inner and outer slope of a NFW profile. Here we test whether our approach can also recover the true potential parameters if the stream is evolving in a dark matter halo with a cored density profile. Cored profiles are particularly interesting for dark matter physics, since they are expected outcomes in self-interacting dark matter (Spergel and Steinhardt, 2000; Bullock and Boylan-Kolchin, 2017). As the outer stream was not as sensitive to the inner slope ($\gamma$) of the potential (see Fig. 2), we attempt to recover $\gamma$ with the inner stream. To run our test, we simulate the inner stream in a cored potential with $\gamma=0.2$ and $\beta=3$.
We rerun our sampler for the inner stream over the same 10 free parameters as described in §IV, where we only fix the on-sky projected position of the progenitor ($x,z$), as well as the stellar disk mass, scale length, and scale height. In Fig. 9, left, we present the best fit stream (black) and the control points used for our sampler (red). In the right panel, we show the recovered inner and outer slope parameters (black lines are the input values), which are both recovered within the $68\%$ region. There is a preference for a non NFW profile at the $2\sigma$ level. We visualize consistent streams from our sampling in Appendix D.
@@ IV.3 Cored density profiles
V Discussion
In this section, we discuss how our method compares to other approaches for modeling extragalactic tidal streams (§V.1), and discuss the limitations of our method (§V.2).
V.1 Comparison to Prior Work
@@
Walder et al. (2025) developed a method to fit orbits to extragalactic stellar streams in a power-law potential of the form $\Phi\propto r^{-\gamma}$. They fit orbits to mock-streams launched from apocenter as a function of different inclination angles. We have reproduced their findings that it is possible to infer the inner slope of the potential (or density) from the turning points of a stream’s track. We have improved upon their method in two main regards. First, we do not fit orbits to the stream’s track, but use a full generative stream model to compare to the observations. Because streams can be significantly misaligned with their progenitor’s orbit, orbit fitting can lead to bias (Sanders and Binney, 2013). Additionally, our approach includes stream width and length information as constraints on the potential. We find that in combination, these can be used to set mass bounds on the halo. Second, our model has more flexibility to model the underlying stream and dark matter halo. In particular, we model the progenitor’s mass and age, while also allowing for variations in an inner and outer halo density slope. We find that multiple streams can be used to stitch together the full radial profile. The disadvantage of our approach is that it is likely slower than the method from Walder et al. (2025), because we compute $1000$ orbits per stream realization. It is worth considering whether a hybrid technique that uses less orbit integrations while also retaining the flexibility we highlight above could lead to further improvements in speed (e.g., the streakline approach Küpper et al. 2012).
V.2 Limitations
@@
For dwarf galaxies, however, much of the stripping occurs near pericenter (see e.g. Bonaca and Price-Whelan, 2025). In Pearson et al. (2022b), they included a Jacobi radius stripping criterion for their dwarf progenitor, which ensured that there was no tidal stripping if the Jacobi radius was much larger than the extent of the progenitor. This in turn led to preferred stripping near pericenter, where the Jacobi radius is smaller. Adding a Jacobi radius criterion to our method would only make it more robust and able to reject more orbits. We can relax the assumption of uniform stripping similar to that of Pearson et al. (2022b) in future work.
Progenitor location: Throughout this paper we fix the present day on-sky position of the progenitor ($x,z$). In observations we will often not have access to the progenitor’s sky position, which introduces some uncertainty to our method (e.g., Martínez-Delgado et al., 2023; Miró-Carretero et al., 2024b). In future work, we can overcome this by marginalizing the progenitor’s location over the stream’s track. This would add a parameter (the angular location of the progenitor along the track), and likely introduce a new set of degeneracies. The method will therefore be more computationally expensive, as we would have to explore many more orbits. However, this additional complexity will most clearly impact our inference of the progenitor properties, but to a lesser extent the radial profile of the host.
@@
Mock observing the streams: In this paper, we applied an “observational” threshold criteria (see §III), but otherwise assumed perfect data. In observations, we likely only observe the densest parts of streams, which will play into the estimated width of each stream from the control points. This could be addressed by adding uncertainties or scatter to our control points representing each stream (see Fig. 4). We do, however, ensure that streams longer than the input data are not disfavored in our sampling.
@@
In halos where two streams are observed, it can be difficult to disentangle whether the two streams are indeed distinct or from the same progenitor (see e.g., Fielder et al. 2025; Bell et al. 2026). Miró-Carretero et al. (2025) found that if we reach a surface brightness limit of 31 mag arcsec-2 stellar tidal streams should be detected around 50% of surveyed hosts, however the occurrence of two streams around one host galaxy might be low. The Roman wide field of view combined with its depth will be ideal for such studies. We also note that streams at smaller galactocentric radii, such as the inner stream, could be more difficult to detect due to a lower surface brightness contrast to the stellar halo. While the detection of multiple dwarf galaxy streams around one host galaxy might be rare and difficult to disentangle, our method can also be used on extragalactic globular cluster streams, predicted to be discovered by Roman (Pearson et al., 2019, 2022a), or on dwarf streams around dwarf galaxies (e.g. Martínez-Delgado et al., 2012; Kado-Fong et al., 2018; Carlin et al., 2019; Fielder et al., 2025).
@@
Time-dependence: Throughout this paper, we have assumed that the host potential and progenitor mass are static. From the Milky Way, we know that time dependence induced from accreted satellites (e.g., Erkal et al., 2019; Shipp et al., 2021; Vasiliev et al., 2021; Arora et al., 2022; Dillamore et al., 2022; Lilleengen et al., 2023; Nibauer et al., 2024; Brooks et al., 2025) and from the overall mass growth of host halos (Buist and Helmi, 2015) can affect the orbits and morphologies of stellar streams. However, depending on the location of the stream in its host halo, the recent merger history, and the dynamical age of the streams, some streams may be less sensitive to time dependence (Nibauer et al., 2022; Brooks et al., 2025). Buist and Helmi (2015) showed that for a smooth halo growth that well reproduced the average mass accretion history of the Milky-Way-sized halos, time evolution led to an angular difference in apocenters at the level of $\approx 10~\deg$, but only for sufficiently long streams which wrapped around their host more than once. In an upcoming work, we test our method on cosmologically evolved halos, where we can determine the bias induced by realistic time-dependence.
[…]
Statistical Assumptions: Our method employs an average likelihood function rather than the true likelihood. This produces better sampling efficiency and is related to the sampling method of likelihood tempering with a constant temperature. To derive significance regions that enclose 68 and 95% of the true posterior mass (rather than the tempered density), we calculate contours based on the assumption of normality. We show in Appendix B that this assumption can introduce moderate bias when drawing contours; however, for heavy-tailed distributions, our contours are conservative (i.e., wider than the true posterior). Nevertheless, we have validated that the method produces stream models consistent with the input data within the 68% confidence regions (e.g., Fig. 6).